Suppose T is the translation described by the vector (0,7) and S is the translation described by the vector (0,6). What vector describes the translation T following S? What vector describes the translation S following T?
Answer by TimothyLamb(4379)

vector for translation T following S:
T + S = (0,7) + (0,6)
T + S = (0,13)
---
vector for translation S following T:
S + T = (0,6) + (0,7)
S + T = (0,13)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php